# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with $L_{2}$-mean Hörmander conditions

### v1

@@ 2. $(ii)$
We first give a useful lemma.
Lemma 4.4.
For a fixed $j\in\operatorname{\mathbb{Z}}$, the following estimates hold for any $Q\in\operatorname{\mathcal{Q}}_{n-j}$ when $n\geq 2$:
  1. (i)
    $\frac{1}{|B_{r_{n-j-1}}|}\int_{\operatorname{\mathbb{R}}^{d}}$ $\Big\|\int_{Q}\Big(\int_{B_{r_{n-j-1}}(c(Q))}\left|K_{j}(x,y)-K_{j}(x,c(Q))\right|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x$
    $\quad\quad\quad\quad\quad\quad\quad\lesssim\big((H_{2}^{2}(n-1)+H_{2}^{2}(n)+2^{-2(n-1)})\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\big)^{\frac{1}{2}};$
  2. (ii)
    $\frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}$ $\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x$
    $\quad\quad\quad\quad\quad\quad\lesssim\left((H_{2}^{2}(n)+H_{2}^{2}(n+1)+2^{-2n})\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\right)^{\frac{1}{2}},$
where $r_{n-j}$ is defined in Theorem 2.4.
Proof.
We first prove $(ii)$. For $n\geq 2$ and a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$, using the same argument as in Lemma 4.3, we can find constants $D_{1,d}$ and $D_{2,d}$ such that the function
$x\mapsto\int_{Q}\Big({\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z}\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y$
is supported on the set
$E_{j}^{Q}:=\big\{x\in\operatorname{\mathbb{R}}^{d}:D_{1,d}2^{j}\leq\text{dist}(x,Q)\leq D_{2,d}2^{j}\big\},$
where $\text{dist}(x,Q):=\inf\left\{|x-y|:y\in Q\right\}$. Then, by the Hölder inequality and the Cauchy-Schwarz inequality, the left-hand side of the inequality $(ii)$ is bounded by $B_{1}\cdot B_{2}$, where $B_{1}$ is defined as
$\frac{1}{|B_{r_{n-j}}|}\bigg(\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x\bigg)^{\frac{1}{2}}$
and $B_{2}$ is defined as
$\Big(\int_{\operatorname{\mathbb{R}}^{d}}\big\|p_{Q}\chi_{E_{j}^{Q}}(x)\big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x\Big)^{\frac{1}{2}}.$
For $B_{1}$, applying the Cauchy-Schwarz inequality to the inner integral for $z$,
$(B_{1})^{2}$ $\leq\int_{Q}\Big(\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y+z)-K_{j}(x,y)|^{2}{\rm d}x{\rm d}z\Big)\tau\big(p_{Q}f(y)p_{Q}\big){\rm d}y$
$\lesssim 2^{-jd}\big(H_{2}^{2}(n)+H_{2}^{2}(n+1)+2^{-2n}\big)\phi(fp_{Q}\chi_{Q}),$
where the last inequality follows from Lemma 4.3 with $q=2$. For $B_{2}$, the construction of $E_{j}^{Q}$ implies that
$(B_{2})^{2}\leq 2^{jd}\tau(p_{Q}).$
Combining the estimates for $B_{1}$ and $B_{2}$, we conclude that
$\frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{\frac{1}{2}(\operatorname{\mathcal{M}})}}^{\frac{1}{2}}{\rm d}x$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\lesssim\left((H_{2}^{2}(n)+H_{2}^{2}(n+1)+2^{-2n})\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\right)^{\frac{1}{2}},$
which completes the proof of $(ii)$.
For $(i)$, we note that there also exist two constants $D_{1,d}^{\prime}$, $D_{2,d}^{\prime}$ such that the support of the function
$x\mapsto\int_{Q}\Big(\int_{B_{r_{n-j-1}(c(Q))}}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u$
is contained in the set
$\big\{x\in\operatorname{\mathbb{R}}^{d}:D_{1,d}^{\prime}2^{j}\leq|x-c(Q)|\leq D_{2,d}^{\prime}2^{j}\big\}.$
The remaining argument parallels that for $(ii)$ and is therefore omitted. ∎
Proof of Lemma 4.2.
@@
Combining this with Remark 2.8, it suffices to prove
@@
We start to estimate the left-hand side. The definition of $\widetilde{b}_{\text{off},n-j}$ in Theorem 2.4 and the cancellation condition in Lemma 2.6 $(iv)$ give:
$\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\quad\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q))\big)\widetilde{b}_{\text{off},n-j}^{Q}(y){\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x.$
Next, recall the definition of $\widetilde{b}_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{r_{n}}$, where $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Accordingly, we decompose $\sum\limits_{n\geq 2}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ into four terms:
@@
for $i=1,2,3,4$, with
$\widetilde{b}_{\text{off},n-j}^{Q,1}:=p_{Q}\big((f\chi_{Q})*h_{r_{n-j}}\big)q_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,2}:=p_{Q}\big((f_{Q}\chi_{Q})*h_{r_{n-j}}\big)q_{Q},$
$\widetilde{b}_{\text{off},n-j}^{Q,3}:=q_{Q}\big((f\chi_{Q})*h_{r_{n-j}}\big)p_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,4}:=q_{Q}\big((f_{Q}\chi_{Q})*h_{r_{n-j}}\big)p_{Q}.$
@@
To estimate $F_{1}$. We first observe that by the supports of $h_{r_{n-j}}$ and $\chi_{Q}$, the noncommutative $L_{1}$-norm in $F_{1}$ equals
$\Big\|\int_{Q}\Big(\int_{|u-y|\leq r_{n-j}}(K_{j}(x,y)-K_{j}(x,c(Q)))h_{r_{n-j}}(y-u){\rm d}y\Big)p_{Q}f(u)q_{Q}{\rm d}u\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
By using Lemma A.3 (in Appendix A), the above estimate is bounded by a product
$\Big\|\Big(\int_{Q}\Big(\int_{B_{r_{n-j}}(u)}\left|K_{i}(x,y)-K_{i}(x,c(Q))\right||h_{r_{n-j}}(y-u)|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big)^{{}^{\frac{1}{2}}}\Big\|_{L_{1}(\operatorname{\mathcal{M}})}$
$\cdot\Big\|\Big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}.$
For the second term, an application of (2.1) shows that
$\Big\|\Big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}\lesssim(|Q|\lambda)^{\frac{1}{2}}.$
@@
$F_{1}\lesssim$ $\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}(|Q|\lambda)^{\frac{1}{2}}$
(4.8) $\frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}(u)}\left|K_{i}(x,y)-K_{i}(x,c(Q))\right|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x.$
The same argument as in (4.2.1) allows us to majorize (4.8) by
$\frac{1}{|B_{r_{n-j-1}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j-1}}(c(Q))}\left|K_{j}(x,y)-K_{j}(x,c(Q))\right|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x.$
By Lemma 4.4 $(i)$ and the $L_{2}$-mean Hörmander condition (1.7), we obtain
$F_{1}$ $\lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\left(\big(H_{2}^{2}(n-1)+H_{2}^{2}(n)+2^{-2(n-1)}\big)\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\right)^{\frac{1}{2}}(|Q|\lambda)^{\frac{1}{2}}$
$\lesssim\sum_{n\geq 2}\left(2H_{2}(n-1)+2^{-n+1}\right)\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}$
$\lesssim\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}.$
Finally, applying the Cauchy-Schwarz inequality twice, we estimate $F_{1}$ as follows:
$F_{1}\lesssim\Big(\sum_{m=1}^{\infty}\lambda\phi(p_{m})\Big)^{\frac{1}{2}}\Big(\sum_{m=1}^{\infty}\phi(fp_{m})\Big)^{\frac{1}{2}}=\big(\lambda\phi(1-q)\big)^{\frac{1}{2}}\big(\phi(f(1-q))\big)^{\frac{1}{2}}$ $\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
$(ii)$ We now consider $b_{\text{off}}^{0}$. Combining (4.7) with Remark 2.8, it suffices to prove
@@
First, a direct calculation gives
$\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)b_{\text{off},n-j}^{0}(y){\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x.$
As in the proof of Lemma 4.1 $(ii)$, we bound the noncommutative $L_{1}$-norm on the right-hand side by
(4.9) $\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\Big\|\int_{Q}\int_{B_{r_{n-j}}}\big(K_{j}(x,y)-K_{j}(x,y+z)\big)h_{r_{n-j}}(z)b^{Q}_{\text{off},n-j}(y){\rm d}z{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
Now, recall the definition of $b_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Substituting this into (4.9), we therefore decompose $\sum\limits_{n\geq 2}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ into four terms:

### v2

@@ 2. $(ii)$
Together with Remark 2.7, it suffices to prove
@@
For a fixed $x\in\operatorname{\mathbb{R}}^{d}$, the definition of $\widetilde{b}_{\text{off},n-j}$ in Theorem 2.4 and the cancellation condition in Lemma 2.5 $(iv)$ give
$\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}(x)=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q))\big)\widetilde{b}_{\text{off},n-j}^{Q}(y){\rm d}y.$
Next, recall the definition of $\widetilde{b}_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)}$, where $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Accordingly, we can majorize $\sum\limits_{n\geq 4}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ by the sum of the following four terms:
@@
for $i=1,2,3,4$, where
$\widetilde{b}_{\text{off},n-j}^{Q,1}:=p_{Q}\big((f\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)q_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,2}:=p_{Q}\big((f_{Q}\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)q_{Q},$
$\widetilde{b}_{\text{off},n-j}^{Q,3}:=q_{Q}\big((f\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)p_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,4}:=q_{Q}\big((f_{Q}\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)p_{Q}.$
@@
Define
$\mathcal{L}_{j}(x,u):=\int_{|u-y|\leq\frac{\sqrt{d}}{2}\ell(Q)}(K_{j}(x,y)-K_{j}(x,c(Q)))h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}y,$
and decompose it into positive and negative parts:
$\mathcal{L}_{j}(x,u)=\mathcal{L}_{j}^{+}(x,u)-\mathcal{L}_{j}^{-}(x,u).$
Hence, for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$, using the supports of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$ and $\chi_{Q}$, the noncommutative $L_{1}$-norm inside $F_{1}$ equals
$\Big\|\int_{Q}\mathcal{L}_{j}(x,u)p_{Q}f(u)q_{Q}{\rm d}u\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
By the triangle inequality, it is enough to treat the positive part, i.e.
(4.10) $\Big\|\int_{Q}\mathcal{L}_{j}^{+}(x,u)p_{Q}f(u)q_{Q}{\rm d}u\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
Let $\widehat{Q}$ be the unique father cube of $Q$ in $\operatorname{\mathcal{Q}}_{n-j-1}$. As in the proof of Lemma 4.1 $(i)$, when $u\in Q$, we can estimate $\mathcal{L}_{j}^{+}(x,u)$ as follows:
$|\mathcal{L}_{j}^{+}(x,u)|$ $\lesssim\frac{1}{|Q|}\int_{|u-y|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y$
$\lesssim\frac{1}{|\widehat{Q}|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(\widehat{Q})}|K_{j}(x,z+c(Q))-K_{j}(x,c(Q))|{\rm d}z$
(4.11) $\leq\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y).$
By Lemma A.4 in Appendix A.3, there exists a contraction $m(x)$ such that
$\int_{Q}\mathcal{L}_{j}^{+}(x,u)p_{Q}f(u)q_{Q}{\rm d}u$
$\quad=\Big(\int_{Q}\mathcal{L}^{+}_{j}(x,u)p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}m(x)\Big(\int_{Q}\mathcal{L}_{j}^{+}(x,u)q_{Q}f(u)q_{Q}d{\rm d}u\Big)^{\frac{1}{2}}$
$\quad=\Big(\int_{Q}\mathcal{L}^{+}_{j}(x,u)p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}p_{Q}m(x)\Big(\int_{Q}\mathcal{L}^{+}_{j}(x,u)q_{Q}f(u)q_{Q}d{\rm d}u\Big)^{\frac{1}{2}}.$
Here we used that $(p_{Q}Xp_{Q})^{\frac{1}{2}}=(p_{Q}Xp_{Q})^{\frac{1}{2}}p_{Q}$ when $X$ is positive. Together it with the Hölder inequality and (4.2.2), we can bound (4.10) by
$\Big\|\Big(\int_{Q}\mathcal{L}_{j}^{+}(x,u)p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}\Big\|p_{Q}m(x)\Big(\int_{Q}\mathcal{L}_{j}^{+}(x,u)q_{Q}f(u)q_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}$
$\lesssim\Big(\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y)\Big)\tau\Big(\int_{Q}p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}\tau\Big(p_{Q}m(x)\big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\big)m(x)^{*}p_{Q}\Big)^{\frac{1}{2}}.$
On the one hand, it is easy to see
$\tau\Big(\int_{Q}p_{Q}f(u)p_{Q}{\rm d}u\Big)\leq\phi\big(fp_{Q}\chi_{Q}\big).$
On the other hand, by the definition of $f_{Q}$, (2.1) and $|m(x)|^{2}\leq 1$,
$\tau\Big(p_{Q}m(x)\big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\big)m(x)^{*}p_{Q}\Big)$ $=\tau\big(|Q|p_{Q}m(x)q_{Q}f_{Q}q_{Q}m(x)^{*}p_{Q}\big)$
$\leq\lambda|Q|\tau\big(p_{Q}m(x)q_{Q}m(x)^{*}p_{Q}\big)$
$\leq\lambda|Q|\tau(p_{Q}).$
@@
[…]
This estimate, combined with Lemma 4.3 and the $L_{1}$-integral condition (1.6), implies
$F_{1}$ $\lesssim\sum_{n\geq 4}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\Big(\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y){\rm d}x\Big)\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}$
$\lesssim\sum_{n\geq 4}\Big(\mathcal{H}(n-3)+\mathcal{H}(n-3)+\mathcal{H}(n-1)+\mathcal{H}(n)+2^{-n+1}\Big)$
$\quad\quad\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}$
$\lesssim\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}.$
Finally, we prove
(4.12) $\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}\lesssim\|f\|_{L_{1}(\operatorname{\mathcal{N}})},$
which gives the desired bound for $F_{1}$. Indeed, by applying the Cauchy-Schwarz inequality twice, we estimate the left-hand side of (4.12) as
$\sum_{n\in\operatorname{\mathbb{Z}}}\Big(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\phi(fp_{Q}\chi_{Q})\Big)^{\frac{1}{2}}\Big(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\tau(p_{Q})(|Q|\lambda\Big)^{\frac{1}{2}}$
$\quad\leq\Big(\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\phi(fp_{Q}\chi_{Q})\Big)^{\frac{1}{2}}\Big(\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\tau(p_{Q})(|Q|\lambda)\Big)^{\frac{1}{2}}$
$\quad=\Big(\sum_{n\geq 1}\phi(fp_{n})\Big)^{\frac{1}{2}}\Big(\sum_{n\geq 1}\lambda\phi(p_{n})\Big)^{\frac{1}{2}}.$
Then, it follows from (2.2) that $\sum\limits_{n\geq 1}\phi(p_{n})=\phi(1-q)$ and $\sum\limits_{n\geq 1}\phi(fp_{n})=\phi\big(f(1-q)\big)$. Combined with Lemma 2.1 $(iv)$, this yields
$\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}\leq\Big(\lambda\phi(1-q)\Big)^{\frac{1}{2}}\Big(\phi(f(1-q))\Big)^{\frac{1}{2}}\lesssim\|f\|_{L_{1}(\operatorname{\mathcal{N}})},$
which proves (4.12) and completes the estimate for $F_{1}$.
$(ii)$ We now consider $b_{\text{off}}^{0}$. By (4.9) and Remark 2.7, it suffices to prove
@@
For a fixed $x\in\operatorname{\mathbb{R}}^{d}$, from the definition of $b_{\text{off},n-j}^{0}$ in Theorem 2.4, a direct calculation implies that $\mathcal{T}_{j}b_{\text{off},n-j}^{0}(x)$ equals to
$\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)(b_{\text{off},n}^{Q}-\widetilde{b}_{\text{off},n}^{Q})(y){\rm d}y$
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\big(K_{j}(x,y)-K_{j}(x,y+z)\big)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)b^{Q}_{\text{off},n-j}(y){\rm d}z{\rm d}y.$
Recall the definition of $b_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Again, $\sum\limits_{n\geq 4}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ can be majorized by the sum of the following four terms: